Find the derivative of the fun

tion

f(x) = 15 − x

6x + 1 +

x

1 + x

.

[1℄ − √

9x+1

6x+1 +

1

(1+x)

2

[2℄ − √

3x

6x+1 +

6x + 1 + 1

(1+x)

2

[3℄ √

3x

6x+1 + 1

(1+x)

2

[4℄ √

9x+1

6x+1 +

1−2x

(1+x)

2


1
Expert's answer
2021-06-17T15:42:18-0400

Solution.

"f(x)=15-x\sqrt{6x+1}+x(1+x)"

Then"f'(x)=-\sqrt{6x+1}-\frac{6x}{2\sqrt{6x+1}}+1+x+x=\newline
-\sqrt{6x+1}-\frac{3x}{\sqrt{6x+1}}+1+2x."